OpenStudy (anonymous):

Simplify (x – 3)(x2 + 7x – 8)

OpenStudy (anonymous):

x^3 + 4x^2 - 29x +24

OpenStudy (anonymous):

Just go along the brackets and multiply each term of the first by every term of the other, then group like terms and simplify: (x*x^2 + x*7x - 8x) - (3*x^2 + 3*7x - 3*8) = x^3 + 7x^2 - 8x - 3x^2 - 21x + 24 = x^3 + 4x^2 - 29x + 24

OpenStudy (anonymous):

my answer same like fewscrewmissing